Find P(A or B or C) for the given probabilities.
Answer #1

P(A U B U C) = P(A) + P(B) + P(C) - P(A^B) - P(B^C) - P(C^A) + P(A^B^C)
P(A U B U C) = 0.35 + 0.23 + 0.18 - 0.13 - 0.03 - 0.07 + 0.01 = 0.54